# On Minimax Estimation of Parameters in Softmax-Contaminated Mixture of Experts

### v1

@@ 4 Numerical Experiments
Training procedure. We conduct $40$ experiments and, for each of them, consider 20 different sample sizes $n$, ranging from $10^{3}$ to $10^{5}$. In computing the MLEs, the initialization is set relatively close to the true parameter values to mitigate potential optimization instabilities. We use an EM algorithm (Jordan-1994, ) to compute the MLE, employing an off-the-shelf BFGS optimizer for the M-step due to the absence of a universal closed-form solution. All the numerical experiments are performed on a MacBook Air with an Apple M4 chip.
@@
(a) Figure 1 displays the results for Theorem 1. We observe that the convergence rates of $(\widehat{\beta}_{n},\widehat{\tau}_{n},\widehat{\eta}_{n},\widehat{\nu}_{n})$ are $\mathcal{O}(n^{-0.45}),\mathcal{O}(n^{-0.52}),\mathcal{O}(n^{-0.50}),\mathcal{O}(n^{-0.54})$, respectively. All the MLEs share similar rates of approximately $\mathcal{O}(n^{-1/2})$, in agreement with the theoretical rates predicted by Theorem 1.
@@ A.1 Proof of Theorem 1
$\lim_{\varepsilon\to 0}\inf_{G,G_{*}}\left\{\frac{\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{D_{1}(G,G_{*})}:D_{1}(G,\overline{G})\vee D_{1}(G_{*},\overline{G})\leq\varepsilon\right\}>0.$
Using the argument with Fatou’s lemma as in Theorem 3.1, Ho-Nguyen-EJS-16 , it is sufficient to show that
@@
In this proof, we will take into account only the most challenging setting of $(\beta_{n},\eta_{n},\nu_{n})$ and $(\beta_{n}^{*},\eta^{*}_{n},\nu^{*}_{n})$ when they converge to the same limit point $(\beta^{\prime},\eta^{\prime},\nu^{\prime})$, where $(\beta^{\prime},\eta^{\prime},\nu^{\prime})$ is not necessarily equal to $(\bar{\beta},\bar{\eta},\bar{\nu})$
@@
$\lim_{n\to\infty}\frac{\mathbb{E}_{X}[V(p_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]}{m_{n}D_{1n}}\geq\int\liminf_{n\to\infty}\frac{|p_{G_{n}}(Y|X)-p_{G_{*}}(Y|X)|}{2m_{n}D_{1n}}d(X,Y)$ (18)
@@
$\mathcal{W}_{1}$ :=$:=\left\{X^{\ell_{1}}\exp((\beta^{*})^{\top}X)\dfrac{\partial^{\alpha_{2}}h}{\partial\eta^{\alpha_{2}}}(X,\eta^{*})\frac{\partial^{\ell_{2}}f}{\partial h^{\ell_{2}}}\left(Y|h(X,\eta^{*}),\nu^{*}\right):0\leq\ell_{2}\leq 2\right\}$
@@
$c_{0}\frac{\partial^{0}f}{\partial h^{0}}+c_{1}\frac{\partial^{1}f}{\partial h}+c_{2}\frac{\partial^{2}f}{\partial h^{2}}+dp_{G_{*}}(Y|X)=0\quad\text{for almost every }Y.$
@@
[…]
@@ A.2 Proof of Theorem 2
Therefore, we will apply the modified Le Cam method for nonsymmetric loss, as outlined in Lemma C.1 of gadat2020parameter , to handle this distance. For $f$ satisfies all assumptions in Theorem 2, based on the Taylor expansion, we have the following results:
@@
[…]
@@
$\mathbb{E}_{X}[h(p_{G_{*}}(\cdot|X),p_{G^{\prime}_{*}}](\cdot|X))\leq C_{0}\epsilon^{r}.$ (20)
Now we will denote $p^{n}_{G_{*}}$ as the density of the $n$-i.i.d. sample $(X_{1},Y_{1}),\cdots,(X_{n},Y_{n})$. Lemma C.1 in gadat2020parameter tells us that
$\inf_{\overline{G}_{n}\in\Xi}\sup_{G\in\Xi}\mathbb{E}_{p_{G}}\Big{(}\exp^{2}(\tau)\|(\overline{\beta}_{n},\overline{\eta}_{n},\overline{\nu}_{n})-(\beta,\eta,\nu)\|^{2}\Big{)}$ $\geq\frac{\epsilon^{2}}{2}\Big{(}1-\mathbb{E}_{X}[V(p^{n}_{G_{*}}(\cdot|X),p^{n}_{G^{\prime}_{*}}(\cdot|X))]\Big{)}$
@@
$p_{S_{j,n}}(Y|X)$ $=\frac{1}{1+\exp(\beta_{j,n}^{\top}X+\tau_{n})}f_{0}(Y|h_{0}(X,\eta_{0}),\nu_{0})+\frac{\exp(\beta_{j,n}^{\top}X+\tau_{n})}{1+\exp(\beta_{j,n}^{\top}X+\tau_{n})}f(Y|h(X,\eta_{j,n}),\nu_{j,n}),$
@@
$\mathbb{E}_{X}$ $[h^{2}(p_{S_{1,n}}(\cdot|X),p_{S_{2,n}}(\cdot|X))]\leq C\int\left(\frac{p_{S_{1,n}}(Y|X)-p_{S_{2,n}}(Y|X)}{p_{S_{2,n}}(Y|X)}\right)^{2}d(X,Y)$
@@
$=\sum_{|\alpha|=1}\frac{(\beta_{1,n}-\beta_{2,n})^{\alpha_{1}}(\eta_{1,n}-\eta_{2,n})^{\alpha_{2}}(\nu_{1,n}-\nu_{2,n})^{\alpha_{3}}}{\alpha_{1}!\alpha_{2}!\alpha_{3}!}X^{\alpha_{1}}\exp(\beta_{2,n}^{\top}X)\frac{\partial^{|\alpha_{2}|+\alpha_{3}}f}{\partial\eta^{\alpha_{2}}\partial\nu^{\alpha_{3}}}(Y|h(X,\eta_{2,n}),\nu_{2,n})$
$\quad+\sum_{|\alpha|=1}\frac{(\beta_{1,n}-\beta_{2,n})^{\alpha_{1}}(\eta_{1,n}-\eta_{2,n})^{\alpha_{2}}(\nu_{1,n}-\nu_{2,n})^{\alpha_{3}}}{\alpha_{1}!\alpha_{2}!\alpha_{3}!}\int_{0}^{1}X^{\alpha_{1}}\exp\left((\beta_{2,n}+t(\beta_{1,n}-\beta_{2,n}))^{\top}X\right)$
$\hskip 142.26378pt\cdot\frac{\partial^{|\alpha_{2}|+\alpha_{3}}f}{\partial\eta^{\alpha_{2}}\partial\nu^{\alpha_{3}}}\left(Y|h(X,\eta_{2,n}+t(\eta_{1,n}-\eta_{2,n})),\nu_{2,n}+t(\nu_{1,n}-\nu_{2,n})\right)dt.$
@@
$\frac{\mathbb{E}_{X}[h^{2}(p_{S_{1,n}}(\cdot|X),p_{S_{2,n}}(\cdot|X))]}{d_{1}^{2r}(S_{1,n},S_{2,n})}\to 0.$
@@
$p_{S^{\prime}_{1,n}}(Y|X)-p_{S^{\prime}_{2,n}}(Y|X)=\frac{e^{\beta_{n}^{\top}X}\left(e^{\tau_{2,n}}-e^{\tau_{1,n}}\right)}{\left(1+e^{\beta_{n}^{\top}X+\tau_{1,n}}\right)\left(1+e^{\beta_{n}^{\top}X+\tau_{2,n}}\right)}\cdot\left[f(Y|h(X,\eta_{n}),\nu_{n})-f_{0}(Y|h_{0}(X,\eta_{0}),\nu_{0})\right].$
@@
$\mathbb{E}_{X}[h^{2}(p_{S^{\prime}_{1,n}}(\cdot|X),p_{S^{\prime}_{2,n}}(\cdot|X))]\leq C\int\left(\frac{p_{S^{\prime}_{1,n}}(Y|X)-p_{S^{\prime}_{2,n}}(Y|X)}{p_{S^{\prime}_{2,n}}(Y|X)}\right)^{2}d(X,Y).$
@@
$\mathbb{E}_{X}[h^{2}(p_{S^{\prime}_{1,n}}(\cdot|X),p_{S^{\prime}_{2,n}}(\cdot|X))]\leq C^{\prime}\left(e^{\tau_{1,n}}-e^{\tau_{2,n}}\right)^{2}.$
@@
$\frac{\mathbb{E}_{X}[h^{2}(p_{S^{\prime}_{1,n}}(\cdot|X),p_{S^{\prime}_{2,n}}(\cdot|X))]}{d_{2}((S^{\prime}_{1,n},S^{\prime}_{2,n}))^{r}}\leq\frac{C^{\prime}\left|e^{\tau_{1,n}}-e^{\tau_{2,n}}\right|^{2}}{\left|e^{\tau_{1,n}}-e^{\tau_{2,n}}\right|^{2r}}=C^{\prime}\left|e^{\tau_{1,n}}-e^{\tau_{2,n}}\right|^{2(1-r)}\to 0$
@@
$\frac{\mathbb{E}_{X}[h^{2}(p_{S^{\prime}_{1,n}}(\cdot|X),p_{S^{\prime}_{2,n}}(\cdot|X))]}{d_{2}^{r}(S^{\prime}_{1,n},S^{\prime}_{2,n})}\to 0,$
@@ A.3 Proof of Theorem 3
${\text{I}}_{n,2,0}(X)=\exp(\tau_{n})\Big{[}\frac{1}{2}(\Delta\nu_{n}-\Delta\nu^{*}_{n})+\sum_{1\leq u,v\leq q}\frac{(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(v)}}{1+\mathbf{1}_{u=v}}\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\frac{\partial h}{\partial\eta^{(v)}}(X,\eta^{*}_{n})\Big{]},$
@@
${\text{II}}_{n,1,0}(X)=\exp(\tau_{n})\Big{[}\sum_{u=1}^{q}\{(-\Delta\eta^{*}_{n})^{(u)}\}\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})+\sum_{1\leq u,v\leq q}\frac{(-\Delta\eta^{*}_{n})^{(u)}(-\Delta\eta^{*}_{n})^{(v)}}{1+\mathbf{1}_{u=v}}\frac{\partial^{2}h}{\partial\eta^{(u)}\partial\eta^{(v)}}(X,\eta^{*}_{n})\Big{]},$
@@
${\text{III}}_{n,1,0}(X)=\exp(\tau^{*}_{n})\Big{[}\sum_{u=1}^{q}\{(-\Delta\eta^{*}_{n})^{(u)}\}\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})+\sum_{1\leq u,v\leq q}\frac{(-\Delta\eta^{*}_{n})^{(u)}(-\Delta\eta^{*}_{n})^{(v)}}{1+\mathbf{1}_{u=v}}\frac{\partial^{2}h}{\partial\eta^{(u)}\partial\eta^{(v)}}(X,\eta^{*}_{n})\Big{]},$
@@
${\text{IV}}_{n,2,0}(X)=\exp(\tau_{n}^{*}+\tau_{n})\Big{[}\frac{1}{2}(\Delta\nu_{n}-\Delta\nu^{*}_{n})+\sum_{1\leq u,v\leq q}\frac{(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(v)}}{1+\mathbf{1}_{u=v}}\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\frac{\partial h}{\partial\eta^{(v)}}(X,\eta^{*}_{n})\Big{]},$
@@
$W_{n}$ $=\sum_{\gamma=0}^{4}\Bigg{[}\left({\text{I}}_{n,\gamma,0}(X)+{\text{II}}_{n,\gamma,0}(X)+{\text{III}}_{n,\gamma,0}(X)\right)+\sum_{\zeta=1}^{2}\left({\text{I}}_{n,\gamma,\zeta}(X)+{\text{II}}_{n,\gamma,\zeta}(X)\right)X^{\zeta}+{\text{IV}}_{n,\gamma,0}(X)\exp((\beta_{n})^{\top}X)\Bigg{]}$
@@
$\mathcal{L}_{0,1}$ =$=\left\{Xf(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
$\mathcal{L}_{0,2}$ =$=\left\{XX^{\top}f(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
$\mathcal{L}_{1,1}$ =$=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})X\frac{\partial f}{\partial h}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[q]\right\}$
$\mathcal{L}_{2,1}$ =$=\left\{X\frac{\partial^{2}f}{\partial h^{2}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
$\mathcal{L}_{1,0}$ =$=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\frac{\partial f}{\partial h}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[d]\right\}$
@@
$\mathcal{L}_{2,0}$ =$=\left\{\frac{\partial^{2}f}{\partial h^{2}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
@@
$\mathcal{L}_{3,0}$ =$=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\frac{\partial^{3}f}{\partial h^{3}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[d]\right\}$
$\mathcal{L}_{4,0}$ =$=\left\{\frac{\partial^{4}f}{\partial h^{4}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\},$
@@
$\mathcal{K}_{1}$ =$=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\exp((\beta_{n})^{\top}X)\frac{\partial f}{\partial h}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[d]\right\}$
@@
$\mathcal{K}_{2}$ =$=\left\{\exp((\beta_{n})^{\top}X)\frac{\partial^{2}f}{\partial h^{2}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
@@
$\mathcal{K}_{3}$ =$=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\exp((\beta_{n})^{\top}X)\frac{\partial^{3}f}{\partial h^{3}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[d]\right\},$
$\mathcal{K}_{4}$ =$=\left\{\exp((\beta_{n})^{\top}X)\frac{\partial^{4}f}{\partial h^{4}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}.$

### v2

@@ 4 Numerical Experiments
Training procedure. We conduct $40$ experiments and, for each of them, consider 20 different sample sizes $n$, ranging from $10^{3}$ to $10^{5}$. In computing the MLEs, the initialization is set relatively close to the true parameter values to mitigate potential optimization instabilities. We use an EM algorithm (Jordan-1994) to compute the MLE, employing an off-the-shelf BFGS optimizer for the M-step due to the absence of a universal closed-form solution. All the numerical experiments are performed on a MacBook Air with an Apple M4 chip.
@@
(a) Figure 1 displays the results for Theorem 1. We observe that the convergence rates of $(\widehat{\beta}_{n},\widehat{\tau}_{n},\widehat{\eta}_{n},\widehat{\nu}_{n})$ are $\mathcal{O}(n^{-0.45}),\mathcal{O}(n^{-0.52}),\mathcal{O}(n^{-0.50}),\mathcal{O}(n^{-0.54})$, respectively, aligning with the theoretical rates of order $\mathcal{O}(n^{-1/2})$ in Theorem 1.
@@ A.1 Proof of Theorem 1
$\lim_{\varepsilon\to 0}\inf_{G,G_{*}}\left\{\frac{\mathbb{E}_{X}[d_{V}(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]}{D_{1}(G,G_{*})}:D_{1}(G,\overline{G})\vee D_{1}(G_{*},\overline{G})\leq\varepsilon\right\}>0.$
Using the argument with Fatou’s lemma as in Theorem 3.1, [Ho-Nguyen-EJS-16] , it is sufficient to show that
@@
In this proof, we will take into account only the most challenging setting of $(\beta_{n},\eta_{n},\nu_{n})$ and $(\beta_{n}^{*},\eta^{*}_{n},\nu^{*}_{n})$ when they converge to the same limit point $(\beta^{\prime},\eta^{\prime},\nu^{\prime})$, where $(\beta^{\prime},\eta^{\prime},\nu^{\prime})$ is not necessarily equal to $(\bar{\beta},\bar{\eta},\bar{\nu})$ .
@@
$\lim_{n\to\infty}\frac{\mathbb{E}_{X}[d_{V}(p_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]}{m_{n}D_{1n}}\geq\int\liminf_{n\to\infty}\frac{|p_{G_{n}}(Y|X)-p_{G_{*}}(Y|X)|}{2m_{n}D_{1n}}d(X,Y)$ (18)
@@
$c_{0}\frac{\partial^{0}f}{\partial h^{0}}+c_{1}\frac{\partial^{1}f}{\partial h}+c_{2}\frac{\partial^{2}f}{\partial h^{2}}+dp_{G_{*}}(Y|X)=0,\quad\text{for almost every }Y.$
@@
$\sum_{i=1}^{d}a_{i}\dfrac{\partial h}{\partial\eta^{(i)}}(X,\eta^{*})=0,\ \text{a.s.},\quad\text{or }\nabla_{a}h(X,\eta^{*})=0,\ \text{a.s.},$
@@ A.2 Proof of Theorem 2
Therefore, we will apply the modified Le Cam method for nonsymmetric loss, as outlined in Lemma C.1 of [gadat2020parameter], to handle this distance. For $f$ satisfies all assumptions in Theorem 2, based on the Taylor expansion, we have the following results:
@@
$\text{(i)}~~\lim_{\epsilon\rightarrow 0}\inf_{S_{1},S_{2}}\left\{\displaystyle\frac{\mathbb{E}_{X}[d_{H}\left(p_{S_{1}}(\cdot|X),p_{S_{2}}(\cdot|X)\right)]}{d_{1}^{r}\left(S_{1},S_{2}\right)}:d_{1}\left(S_{1},S_{2}\right)\leq\epsilon\right\}=0,$
$\text{(ii)}~~\lim_{\epsilon\rightarrow 0}\inf_{S^{\prime}_{1},S^{\prime}_{2}}\left\{\displaystyle\frac{\mathbb{E}_{X}[d_{H}\left(p_{S^{\prime}_{1}}(\cdot|X),p_{S^{\prime}_{2}}(\cdot|X)\right)]}{d_{2}^{r}\left(S^{\prime}_{1},S^{\prime}_{2}\right)}:d_{2}\left(S^{\prime}_{1},S^{\prime}_{2}\right)\leq\epsilon\right\}=0.$
@@
$\mathbb{E}_{X}[d_{H}(p_{G_{*}}(\cdot|X),p_{G^{\prime}_{*}}](\cdot|X))\leq C_{0}\epsilon^{r}.$ (20)
Now we will denote $p^{n}_{G_{*}}$ as the density of the $n$-i.i.d. sample $(X_{1},Y_{1}),\cdots,(X_{n},Y_{n})$. Lemma C.1 in [gadat2020parameter] tells us that
$\inf_{\overline{G}_{n}\in\Xi}\sup_{G\in\Xi}\mathbb{E}_{p_{G}}\Big(\exp^{2}(\tau)\|(\overline{\beta}_{n},\overline{\eta}_{n},\overline{\nu}_{n})-(\beta,\eta,\nu)\|^{2}\Big)$ $\geq\frac{\epsilon^{2}}{2}\Big(1-\mathbb{E}_{X}[d_{V}(p^{n}_{G_{*}}(\cdot|X),p^{n}_{G^{\prime}_{*}}(\cdot|X))]\Big)$
@@
$p_{S_{j,n}}(Y|X)$ $=\frac{1}{1+\exp(\beta_{j,n}^{\top}X+\tau_{n})}f_{0}(Y|h_{0}(X,\eta_{0}),\nu_{0})$
$\hskip 113.81102pt+\frac{\exp(\beta_{j,n}^{\top}X+\tau_{n})}{1+\exp(\beta_{j,n}^{\top}X+\tau_{n})}f(Y|h(X,\eta_{j,n}),\nu_{j,n}),$
@@
$\mathbb{E}_{X}$ $[d_{H}^{2}(p_{S_{1,n}}(\cdot|X),p_{S_{2,n}}(\cdot|X))]\leq C\int\left(\frac{p_{S_{1,n}}(Y|X)-p_{S_{2,n}}(Y|X)}{p_{S_{2,n}}(Y|X)}\right)^{2}d(X,Y)$
@@
$=\sum_{|\alpha|=1}\frac{(\beta_{1,n}-\beta_{2,n})^{\alpha_{1}}(\eta_{1,n}-\eta_{2,n})^{\alpha_{2}}(\nu_{1,n}-\nu_{2,n})^{\alpha_{3}}}{\alpha_{1}!\alpha_{2}!\alpha_{3}!}$
$\hskip 142.26378pt\cdot X^{\alpha_{1}}\exp(\beta_{2,n}^{\top}X)\frac{\partial^{|\alpha_{2}|+\alpha_{3}}f}{\partial\eta^{\alpha_{2}}\partial\nu^{\alpha_{3}}}(Y|h(X,\eta_{2,n}),\nu_{2,n})$
$+\sum_{|\alpha|=1}\frac{(\beta_{1,n}-\beta_{2,n})^{\alpha_{1}}(\eta_{1,n}-\eta_{2,n})^{\alpha_{2}}(\nu_{1,n}-\nu_{2,n})^{\alpha_{3}}}{\alpha_{1}!\alpha_{2}!\alpha_{3}!}\int_{0}^{1}X^{\alpha_{1}}\exp\left((\beta_{2,n}+t(\beta_{1,n}-\beta_{2,n}))^{\top}X\right)$
$\hskip 113.81102pt\cdot\frac{\partial^{|\alpha_{2}|+\alpha_{3}}f}{\partial\eta^{\alpha_{2}}\partial\nu^{\alpha_{3}}}\left(Y|h(X,\eta_{2,n}+t(\eta_{1,n}-\eta_{2,n})),\nu_{2,n}+t(\nu_{1,n}-\nu_{2,n})\right)dt.$
@@
$\frac{\mathbb{E}_{X}[d_{H}^{2}(p_{S_{1,n}}(\cdot|X),p_{S_{2,n}}(\cdot|X))]}{d_{1}^{2r}(S_{1,n},S_{2,n})}\to 0.$
@@
$p_{S^{\prime}_{1,n}}(Y|X)-p_{S^{\prime}_{2,n}}(Y|X)$ $=\frac{e^{\beta_{n}^{\top}X}\left(e^{\tau_{2,n}}-e^{\tau_{1,n}}\right)}{\left(1+e^{\beta_{n}^{\top}X+\tau_{1,n}}\right)\left(1+e^{\beta_{n}^{\top}X+\tau_{2,n}}\right)}$
$\hskip 85.35826pt\cdot\left[f(Y|h(X,\eta_{n}),\nu_{n})-f_{0}(Y|h_{0}(X,\eta_{0}),\nu_{0})\right].$
@@
$\mathbb{E}_{X}[d_{H}^{2}(p_{S^{\prime}_{1,n}}(\cdot|X),p_{S^{\prime}_{2,n}}(\cdot|X))]\leq C\int\left(\frac{p_{S^{\prime}_{1,n}}(Y|X)-p_{S^{\prime}_{2,n}}(Y|X)}{p_{S^{\prime}_{2,n}}(Y|X)}\right)^{2}d(X,Y).$
@@
$\mathbb{E}_{X}[d_{H}^{2}(p_{S^{\prime}_{1,n}}(\cdot|X),p_{S^{\prime}_{2,n}}(\cdot|X))]\leq C^{\prime}\left(e^{\tau_{1,n}}-e^{\tau_{2,n}}\right)^{2}.$
@@
$\frac{\mathbb{E}_{X}[d_{H}^{2}(p_{S^{\prime}_{1,n}}(\cdot|X),p_{S^{\prime}_{2,n}}(\cdot|X))]}{d_{2}((S^{\prime}_{1,n},S^{\prime}_{2,n}))^{r}}\leq\frac{C^{\prime}\left|e^{\tau_{1,n}}-e^{\tau_{2,n}}\right|^{2}}{\left|e^{\tau_{1,n}}-e^{\tau_{2,n}}\right|^{2r}}=C^{\prime}\left|e^{\tau_{1,n}}-e^{\tau_{2,n}}\right|^{2(1-r)}\to 0$
@@
$\frac{\mathbb{E}_{X}[d_{H}^{2}(p_{S^{\prime}_{1,n}}(\cdot|X),p_{S^{\prime}_{2,n}}(\cdot|X))]}{d_{2}^{r}(S^{\prime}_{1,n},S^{\prime}_{2,n})}\to 0,$
@@ A.3 Proof of Theorem 3
${\text{I}}_{n,2,0}(X)=\exp(\tau_{n})\Big[\frac{1}{2}(\Delta\nu_{n}-\Delta\nu^{*}_{n})+$
$\hskip 92.47145pt\sum_{1\leq u,v\leq q}\frac{(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(v)}}{1+\mathbf{1}_{u=v}}\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\frac{\partial h}{\partial\eta^{(v)}}(X,\eta^{*}_{n})\Big],$
@@
${\text{II}}_{n,1,0}(X)=\exp(\tau_{n})\Big[\sum_{u=1}^{q}\{(-\Delta\eta^{*}_{n})^{(u)}\}\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})$
$\hskip 136.5733pt+\sum_{1\leq u,v\leq q}\frac{(-\Delta\eta^{*}_{n})^{(u)}(-\Delta\eta^{*}_{n})^{(v)}}{1+\mathbf{1}_{u=v}}\frac{\partial^{2}h}{\partial\eta^{(u)}\partial\eta^{(v)}}(X,\eta^{*}_{n})\Big],$
@@
${\text{III}}_{n,1,0}(X)=\exp(\tau^{*}_{n})\Big[\sum_{u=1}^{q}\{(-\Delta\eta^{*}_{n})^{(u)}\}\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})$
$\hskip 136.5733pt+\sum_{1\leq u,v\leq q}\frac{(-\Delta\eta^{*}_{n})^{(u)}(-\Delta\eta^{*}_{n})^{(v)}}{1+\mathbf{1}_{u=v}}\frac{\partial^{2}h}{\partial\eta^{(u)}\partial\eta^{(v)}}(X,\eta^{*}_{n})\Big],$
@@
${\text{IV}}_{n,2,0}(X)=\exp(\tau_{n}^{*}+\tau_{n})\Big[\frac{1}{2}(\Delta\nu_{n}-\Delta\nu^{*}_{n})$
$\hskip 93.89418pt+\sum_{1\leq u,v\leq q}\frac{(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(v)}}{1+\mathbf{1}_{u=v}}\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\frac{\partial h}{\partial\eta^{(v)}}(X,\eta^{*}_{n})\Big],$
@@
$W_{n}$ $=\sum_{\gamma=0}^{4}\Bigg[\left({\text{I}}_{n,\gamma,0}(X)+{\text{II}}_{n,\gamma,0}(X)+{\text{III}}_{n,\gamma,0}(X)\right)$
$\hskip 51.21504pt+\sum_{\zeta=1}^{2}\left({\text{I}}_{n,\gamma,\zeta}(X)+{\text{II}}_{n,\gamma,\zeta}(X)\right)X^{\zeta}+{\text{IV}}_{n,\gamma,0}(X)\exp((\beta_{n})^{\top}X)\Bigg]$
@@
$\mathcal{L}_{0,1}$ $=\left\{Xf(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
$\mathcal{L}_{0,2}$ $=\left\{XX^{\top}f(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
$\mathcal{L}_{1,1}$ $=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})X\frac{\partial f}{\partial h}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[q]\right\}$
$\mathcal{L}_{2,1}$ $=\left\{X\frac{\partial^{2}f}{\partial h^{2}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
$\mathcal{L}_{1,0}$ $=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\frac{\partial f}{\partial h}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[d]\right\}$
@@
$\mathcal{L}_{2,0}$ $=\left\{\frac{\partial^{2}f}{\partial h^{2}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
@@
$\mathcal{L}_{3,0}$ $=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\frac{\partial^{3}f}{\partial h^{3}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[d]\right\}$
$\mathcal{L}_{4,0}$ $=\left\{\frac{\partial^{4}f}{\partial h^{4}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\},$
@@
$\mathcal{K}_{1}$ $=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\exp((\beta_{n})^{\top}X)\frac{\partial f}{\partial h}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[d]\right\}$
@@
$\mathcal{K}_{2}$ $=\left\{\exp((\beta_{n})^{\top}X)\frac{\partial^{2}f}{\partial h^{2}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}$
@@
$\mathcal{K}_{3}$ $=\left\{\frac{\partial h}{\partial\eta^{(u)}}(X,\eta^{*}_{n})\exp((\beta_{n})^{\top}X)\frac{\partial^{3}f}{\partial h^{3}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X):u\in[d]\right\},$
$\mathcal{K}_{4}$ $=\left\{\exp((\beta_{n})^{\top}X)\frac{\partial^{4}f}{\partial h^{4}}(Y|h(X,\eta_{n}^{*}),\nu_{n}^{*}))\exp((\beta_{n}^{*})^{\top}X)\right\}.$